OpenStudy (anonymous):

The difference of the two angles is 10 degree and the sum of the angles in circular measure is 2. Find the circular measure of each angle.

OpenStudy (mww):

I would recommend you convert degrees into circular measure (radians?) remember that 180 degrees in pi radians. So 10 degrees is (10/180) of pi radians = pi/18 radians. Now set up two simultaneous eqns. Let A and B be the two angles, where A > B. A - B = 10 degrees = pi/18 rad (i) A + B = 2 rad (ii) So (i) + (ii) --> 2A = pi/18 + 2. A = (pi/36 + 1) radians Substitute to find B: B = 2 - A = 2 - (pi/36 + 1) = 2 - pi/36 - 1 = (1 - pi/36) radians